OpenStudy (anonymous):

Can anyone help explain and solve please?? The temperature of a chemical reaction ranges between 40°C and 180°C. The temperature is at its lowest point when t = 0 and completes one cycle over a 12-hour period. What is a sine function that would model this reaction?

OpenStudy (anonymous):

f(x) = 70 sin 12x + 110 f(x) = 110 sin 12x + 70 f(x) = 70 sin pi over 6x + 110 f(x) = 110 sin pi over 6x + 70

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

range if of length \(180-40=140\) and therefor the number out front should be \(70\) ignore 2 and 5

OpenStudy (anonymous):

since the period is \(12\) you can set \(\frac{2\pi}{b}=12\) and solve for \(b\)

OpenStudy (anonymous):

you should get \(b=\frac{\pi}{6}\)

OpenStudy (anonymous):

What about the numbers being added? (in the beginning of each answer) what do they mean/how do I get those numbers? @satellite73

OpenStudy (anonymous):

@thomaster @robtobey @abb0t @iambatman @RadEn

OpenStudy (dumbcow):

\[y = A \sin (Bx + C) + k\] min = k - A max = k + A period = 2pi/B shift is determined by "C" given point is that lowest point occurs at t=0 ..... (0, 40) thus sin(Bx +C) = -1 when x=0 sin(C) = -1 C = -pi/2